# 3rd Generation Partnership Project 5G Fifth Generation

### v1

@@ 10 Gigabit PON
This version of the paper has been submitted for publication to IEEE and may be removed upon request or copyright issues. xApp-Level Conflict Mitigation in O-RAN, a Mobility Driven Energy Saving Case
@@ Abstract
This paper investigates the emerging challenges of conflict detection and mitigation in Open Radio Access Network (O-RAN). Conflicts between xApps can arise that affect network performance and stability due to the disaggregated nature of O-RAN. This work provides a detailed theoretical framework of Extended Application (xApp)-level conflicts, i.e., direct, indirect, and implicit conflicts. Leveraging conflict graphs, we further highlight how conflicts impact Key Performance Indicators and explore strategies for conflict detection using Service Level Agreements and Quality of Service (QoS) thresholds. We evaluate the effectiveness of several mitigation strategies in a simulated environment with Mobility Robustness Optimization (MRO) and Energy Saving (ES) xApps and present experimental results showing comparisons among these strategies. The findings of this research provide significant insights for enhancing O-RAN deployments with flexible and efficient conflict management.
@@ I Introduction
The introduction of O-RAN architecture has promised enhanced flexibility and interoperability in the telecommunications industry. However, this disaggregated approach of integrating components from multiple vendors presents a significant challenge of conflict management. While the traditional Radio Access Network (RAN) relied on single-vendor solutions that come with in-house vendor spesific conflict resolution strategies, O-RAN’s multi-vendor nature needs a robust conflict mitigation strategy to cope with conflicts between different vendors’ components. Conflicts arise when components compete for shared resources or set conflicting configurations that control the same network environment. This can lead to performance degradation, instability, and security vulnerabilities in the network. Therefore, handling these conflicts are now a priority and a major obstcle for wide deployment on O-RAN. The following provides an initial understanding of different types of conflicts in O-RAN based on the state-of-the-art.
I-A Categorization of Conflicts:
Conflicts are categorized into horizontal and vertical conflicts, which are further classified into intra-component and inter-component conflicts [1, 2, 3].
Horizontal conflicts occur between components at the same hierarchical level, for instance, between two xApps (applications for near-real-time control) within a Near Real Time RAN Intelligent Controller (Near-RT-RIC) [2]. It is further classified into intra-component and inter-component conflicts:
    • Indirect Conflicts: These arise when different xApps control distinct parameters that ultimately influence the KPI [1, 6, 8, 2, 9]. For example, one xApp might modify the cell individual offset (CIO) to balance the load across cells, while another adjusts antenna tilts. Both actions indirectly impact the cell boundaries and handover decisions, potentially causing conflicts [10, 8]. Detecting indirect conflicts necessitates post-action analysis, observing how different xApp actions collectively impact the relevant KPI. This is more challenging than detecting direct conflicts as the relationship between parameter changes and KPI impacts might not be immediately apparent [3].
    • Implicit Conflicts: These present the most significant challenge for detection and mitigation [1, 7]. Implicit conflicts occur when the actions of multiple xApps, while individually aligned with their specific objectives, result in an undesirable overall network state [4, 8]. This often involves subtle interactions between xApps with different goals, making the source of the conflict difficult to pinpoint. For example, an xApp trying to maximize quality of service (QoS) for a group of users might interfere with another xApp aiming to minimize handover rates, leading to unexpected performance degradation. Identifying and resolving implicit conflicts often requires sophisticated monitoring and analysis of network behavior over time [1].
Vertical conflicts involve components from different hierarchical levels, such as a conflict between a Near-RT-RIC and a Non-RT RIC [2, 3].
I-A1 Conflict Detection and Mitigation Frameworks:
Adamczyk and Kliks [1] propose a conflict mitigation framework (CMF) within the Near-RT-RIC. The framework focuses on detecting direct, indirect, and implicit conflicts between xApps. Direct conflicts arise from consecutive, contradictory decisions affecting the same parameters. Indirect conflicts occur when xApps impact shared KPIs through different parameters. Implicit conflicts involve xApp actions aligning with individual objectives but contradicting overall network goals. A more detailed description of these conflicts are discussed in Section. II. The framework in [1] utilizes a Conflict Mitigation (CM) component in the Near-RT-RIC to identify and resolve these conflicts. However, current implementations primarily focus on specific xApps, like Mobility Robustness Optimization (MRO) and Mobility Load Balancing (MLB).
I-A2 Team Learning and Cooperative Approaches:
I-A3 QoS-Aware Conflict Mitigation:
I-A4 Motivation and Contribution of this Research:
@@
  • As Open RAN is a new concept, and there is limited knowledge about conflicts within it. The subtle difference between indirect and implicit conflicts makes reliably detecting and evaluating these potential issues challenging. Developing a low-latency detection mechanism is crucial to ensure the reliability and stability of O-RAN deployments.
@@
  • Scalability and Complexity: As O-RAN networks expand in size and complexity, managing conflicts between numerous xApps and Remote Applications from diverse vendors will become increasingly challenging. Scalable and efficient conflict detection and mitigation mechanisms are essential for the large-scale adoption of Open RAN.
@@
The structure of the paper is as follows: Section II provides a theoretical modeling and evaluation of xApp conflicts, Section III presents a discussion on conflict detection and mitigation strategies, Section IV discusses simulation experiments and results with ES and MRO xApps, and the paper concludes in Section V.
II Understanding xApp Conflicts in Open RAN
II-A Theoretical Modeling of Intra-Component Conflicts
We develop a theoretical framework for modeling intra-component conflicts, focusing on the interactions between different xApps within the Near-RT-RIC. This framework uses conflict graphs to represent potential conflict points and their characteristics. Let us consider, there are $n$ number of xApps present in a Near-RT-RIC, represented by $x_{n}\in X$. Additionally, there are $i$ numbers of Input Control Parameters and $j$ number of KPIs, each represented as $p_{i}\in P$ and $k_{j}\in K$, respectively. Each xApp has a set of ICPs denoted by $I_{x_{n}}$ and a set of KPIs represented by $K_{x_{n}}$. Intra-component conflicts are of three types as stated above in Section I-A.
@@ II-A3 Implicit Conflict
III Conflict Detection and Mitigation
Detecting conflicts is one of the most challenging aspects of conflict management in the Near-RT-RIC. It requires a deep understanding of each type of conflict and their distinctions. The comprehensive theoretical analysis presented in Section II and the comparison in Table I provide the foundation for a rule-based, low-latency conflict detection mechanism. xApp-level conflicts are typically parametric and can often be detected by tracing ICPs and monitoring KPI degradation events. Some KPIs are critical to maintain as per SLA agreements, while others are less so.
@@ III Conflict Detection and Mitigation
The Near-RT-RIC database consists of several components, including Recently Changed Parameter (RCP), Parameter Group Definition (PGD), Recently Changed Parameter Group (RCPG), Parameter and KPI Ranges (PKR), Decision Correlated with KPI Degradation (DCKD), and KPI Degradation Occurrences (KDO). When a xApp is deployed in the Near-RT-RIC, the MNO is expected to provide the xApp details, including their ICPs ($I_{x_{n}}$) and KPIs ($K_{x_{n}}$). The PGD component of the database is equipped with Algorithm 1, which estimates the $P_{k}^{G}$ for all $k_{j}\in K_{x_{n}}$. The KDO component is responsible for storing and tracking KPI degradation events. If any SLA-sensitive KPI falls below the predefined threshold, it alerts the CDC of a possible conflict, along with the timestamp ($t_{clock}$) indicating when the event occured.
The RCP component tracks any parameter changes, aiding the CDC in identifying the exact parameter responsible for the conflict at $t_{clock}$. The step-by-step conflict detection process is illustrated in Fig. 4, which explains how a rule-based detection mechanism can identify specific types of conflicts.
@@
When the conflict is detected, the CDC informs the CMC about the specific types of conflict occurring over the specific conflicting ICP. For the analysis of this paper, we adopt the priority conflict mitigation method proposed by Adamczyk et al. [1] and QoS-Aware Conflict Mitigation (QACM) method proposed in [3] to be deployed in the CMC. The priority method simply prioritises one conflicting xApp over the other based on their significance set by the MNO. On the other hand, the QACM considers specific QoS requirements of the conflicting KPIs and suggests an optimal value of the conflicting parameter at which the KPIs are either above their QoS mandates or as closer as possible to it. The following section discusses the simulation setup and results based on these two mitigation methods.
Figure 5: Network layout of the simulated O-RAN.
IV Experiment and Results
For the experimental analysis, we used MATLAB software for simulation, along with its 5G Toolbox and O-RAN 7.2 split [12, 3]. The simulation setup included two Next-generation NodeBs and ten User Equipments, as shown in Fig. 5. We focused on a direct conflict between the MRO xApp and ES xApp over Tx Power (TXP) values. The simulation parameters were: a Reference Signal Received Power (RSRP) threshold of -110 dBm for handover, a frequency of 2.4 GHz, and a simulation time of 10 minutes with a time step of 100 ms. Input control parameters included a default TXP value of 30 dBm, Cell Individual Offset (CIO) of 2 dB, Handover Hysteresis (HYS) of 0.5 dB, Time-To-Trigger (TTT) of 0.1 ms, Radio Electrical Tilt (RET) of 1.5 degrees, and an adjustment interval of 1000 ms. The UEs moved back and forth between the two gNBs with randomly assigned velocities between 0 and 5 m/s. The overall system goal is to minimize link failures and ping-pong handovers for MRO while maximizing energy efficiency for ES. A direct conflict arose when ES set the TXP to 3 dBm for energy savings, followed by MRO setting it to 50 dBm in the next request. We consider 5 conflict mitigation Methods, including a no-mitigation approach (NC) where the TXP was set to the recently requested value; a set back to default (SBD) method where the TXP was reset to its default value; a priority-based method where either ES (P-ES) or MRO (P-MRO) was prioritized; and finaly the QACM method was used to balance the TXP based on QoS thresholds. We repeated the simulation 500 times and the results were presented using box plots for each KPI.
@@ IV Experiment and Results
(c) Number of ping-pong handover.
@@
The Fig. 6(a) compares the energy efficiency across different conflict mitigation methods, including NC (No Conflict Mitigation), SBD (Set Back to Default), P-ES (prioritized ES), P-MRO (prioritized MRO), and QACM. The QACM method shows the highest energy efficiency (in bits/Joule), suggesting it is the most effective approach for maximizing energy savings, followed by P-ES. The NC, SBD and P-MRO methods exhibit lower energy efficiency indicating that prioritizing energy-saving strategies improves this KPI.
@@
The Fig. 6(b) illustrates the impact of different conflict mitigation methods on reducing link failures. The SDB displays higher numbers of link failures followed by the NC and P-ES, implying poorer performance in maintaining stable connections. On the other hand, P-MRO and QACM show significantly fewer link failures with QACM being the most effective one. This indicates that prioritizing mobility and a balanced conflict resolution approach reduces the occurrence of link failures.
@@
Finally, the Fig. 6(c) demonstrates the number of ping-pong handovers across the various methods. NC, SBD and P-ES have the highest frequency of ping-pong handovers, which may negatively affect network stability. In contrast, P-MRO and QACM exhibit significantly fewer ping-pong handovers with QACM providing the best mitigation. This shows that a mobility-optimized and balanced conflict resolution strategy can effectively reduce these undesirable handovers.
V Conclusion

### v2

@@ 10 Gigabit PON
This version of the paper has been accepted for publication to IEEE INFOCOM and may be removed upon request or copyright issues. xApp-Level Conflict Mitigation in O-RAN, a Mobility Driven Energy Saving Case
@@ Abstract
This paper addresses the challenges of conflict detection and mitigation in Open Radio Access Network (RAN) where multiple vendors co-operate the same RAN and share access to control parameters and conflicts between Extended Applications can arise that affect network performance and stability due to the disaggregated nature of Open RAN. This work provides a detailed theoretical framework of xApp-level conflicts, i.e., direct, indirect, and implicit conflicts. Leveraging conflict graphs, we further highlight how conflicts impact Key Performance Indicators and propose a low-latency rule-based strategy for conflict detection using Service Level Agreements and Quality of Service (QoS) thresholds. We evaluate the effectiveness of several mitigation strategies in a simulated environment with Mobility Robustness Optimization (MRO) and Energy Saving (ES) xApps and present experimental results showing comparisons among these strategies. The findings of this research provide significant insights for enhancing O-RAN deployments with flexible and efficient conflict management.
@@ I Introduction
The introduction of Open RAN architecture has promised enhanced flexibility and interoperability in the telecommunications industry. However, this disaggregated approach of integrating components from multiple vendors presents a significant challenge of conflict management. While the traditional RAN relied on single-vendor solutions that come with in-house vendor specific conflict resolution strategies, Open RAN’s multi-vendor nature needs a robust conflict mitigation strategy to cope with conflicts between different vendors’ components. Conflicts arise when components compete for shared resources or set conflicting configurations that control the same network environment. This can lead to performance degradation, instability, and security vulnerabilities in the network. Therefore, handling these conflicts has become a priority and a major obstacle to the widespread deployment of Open RAN.
A key challenge is that the conflict definitions provided by the O-RAN Alliance are not comprehensive, leading to various interpretations of conflicts. Some research contradicts the basic conflict definitions established by the O-RAN Alliance in [6]. For instance, in [5, 13], the definition of implicit conflict aligns with what is typically described as an indirect conflict—where different parameters belonging to separate xApps influence the same KPI. To provide context, the O-RAN Alliance defines that Antenna Tilt (or Radio Electrical Tilt (RET)), associated with Mobility Load Balancing (MLB)/ Capacity and Coverage Optimization (CCO) xApps, affects the handover boundary, while the measurement offset related to MRO xApp has a similar impact. This results in handover-related KPIs being influenced by parameters from distinct xApps, leading to what is described as an indirect conflict. However, in these studies, the definition of indirect conflict has been altered and labeled as implicit conflict that makes their entire detection and mitigation strategy flawed. Similarly, their definition of indirect conflict describes a parameter-to-parameter indirect relationship, which does not align with the O-RAN Alliance’s definition of parameter-to-KPI relationship. Therefore, this paper provides a comprehensive guide to conflict definitions in Open RAN, ensuring alignment with the O-RAN Alliance’s framework.
@@
  • Knowledge Gap: As Open RAN is a new concept, and there is limited knowledge and misconceptions about conflicts within it, therefore, this paper provides a comprehensive discussion on various types of conflicts in Open RAN.
  • Indirect vs Implicit Conflict: The subtle difference between indirect and implicit conflicts makes reliably detecting and evaluating these potential issues challenging. Developing a low-latency detection mechanism is crucial to ensure the reliability and stability of Open RAN deployments.
@@
  • Scalability and Complexity: As Open RAN networks expand in size and complexity, managing conflicts between numerous xApps and Remote Applications from diverse vendors will become increasingly challenging. Scalable and efficient conflict detection and mitigation mechanisms are essential for the large-scale adoption of Open RAN.
@@
  • Studying the definition of conflicts based on state-of-art and O-RAN Alliance guideline in a comprehensive way to clear misconceptions.
@@
The structure of the paper is as follows: Section II provides a comprehensive overview of conflicts in Open RAN, Section III discusses the system model and discusses xApp conflicts with a theoretical model, Section IV presents a discussion on the rule-based conflict detection method and QoS-Aware Conflict Mitigation (QACM) mitigation strategy, Section V discusses simulation experiments and results with ES and MRO xApps, Section VI presents the limitations of this work indicating future research directions, and the paper concludes in Section VII.
II Background Study and State-of-the-Art
This section provides an initial understanding of different types of conflicts in Open RAN based on the state-of-the-art. Conflicts are categorized into vertical and horizontal conflicts, which are further classified into intra-component and inter-component conflicts [1, 2, 3].
Vertical conflicts involve components from different hierarchical levels, such as a conflict between a Near Real Time RAN Intelligent Controller (Near-RT-RIC) and a Non-RT RIC [2, 3]. Horizontal conflicts occur between components at the same hierarchical level, for instance, between two xApps (applications for near-real-time control) within a Near-RT-RIC [2]. It is further classified into intra-component and inter-component conflicts:
    • Indirect Conflicts: These arise when different xApps control distinct parameters that ultimately influence the KPI [1, 6, 8, 2, 9]. For example, MRO xApp might modify the cell individual offset (CIO) to adjust the handover success rate and reduce link failures, while MLB adjusts antenna tilts to balance the traffic loads. Both actions indirectly impact the cell boundaries and handover decisions, potentially causing conflicts. Detecting indirect conflicts necessitates post-action analysis, observing how different xApp actions collectively impact the relevant KPI. This is more challenging than detecting direct conflicts as the relationship between parameter changes and KPI impacts might not be immediately apparent [3].
    • Implicit Conflicts: These present the most significant challenge for detection and mitigation [1, 7]. Implicit conflicts occur when the actions of multiple xApps, while individually aligned with their specific objectives, result in an undesirable overall network state [4, 8]. This often involves subtle interactions between xApps with different goals, making the source of the conflict difficult to pinpoint. For example, a SLA-assuring xApp trying to maximize quality of service (QoS) for a group of users might interfere with another xApp aiming to minimize handover rates, leading to unexpected performance degradation. Identifying and resolving implicit conflicts often requires sophisticated monitoring and analysis of network behavior over time [1].
II-1 Conflict Detection and Mitigation Frameworks:
Adamczyk and Kliks [1] propose a conflict mitigation framework (CMF) within the Near-RT-RIC. The framework focuses on detecting direct, indirect, and implicit conflicts between xApps. Direct conflicts arise from consecutive, contradictory decisions affecting the same parameters. Indirect conflicts occur when xApps impact shared KPIs through different parameters. Implicit conflicts involve xApp actions aligning with individual objectives but contradicting overall network goals. A more detailed description of these conflicts are discussed in Section. III. The framework in [1] utilizes a Conflict Mitigation (CM) component in the Near-RT-RIC to identify and resolve these conflicts. However, current implementations primarily focus on specific xApps, like Mobility Robustness Optimization (MRO) and Mobility Load Balancing (MLB).
II-2 Team Learning and Cooperative Approaches:
II-3 QoS-Aware Conflict Mitigation:
III System Model
We develop a theoretical framework for modeling intra-component conflicts, focusing on the interactions between different xApps within the Near-RT-RIC. This framework uses conflict graphs to represent potential conflict points and their characteristics. Let us consider, there are $n$ number of xApps present in a Near-RT-RIC, represented by $x_{n}\in X$. Additionally, there are $i$ numbers of Input Control Parameters and $j$ number of KPIs, each represented as $p_{i}\in P$ and $k_{j}\in K$, respectively. Each xApp has a set of ICPs denoted by $I_{x_{n}}$ and a set of KPIs represented by $K_{x_{n}}$. Intra-component conflicts are of three types as stated above in Section II.
@@ III-3 Implicit Conflict
IV Conflict Detection and Mitigation
Detecting conflicts is one of the most challenging aspects of conflict management in the Near-RT-RIC. It requires a deep understanding of each type of conflict and their distinctions. The comprehensive theoretical analysis presented in Section III and the comparison in Table I provide the foundation for a rule-based, low-latency conflict detection mechanism. xApp-level conflicts are typically parametric and can often be detected by tracing ICPs and monitoring KPI degradation events. Some KPIs are critical to maintain as per SLA agreements, while others are less so.
@@ IV Conflict Detection and Mitigation
The Near-RT-RIC database consists of several components, including Recently Changed Parameter (RCP), Parameter Group Definition (PGD), Recently Changed Parameter Group (RCPG), Parameter and KPI Ranges (PKR), Decision Correlated with KPI Degradation (DCKD), and KPI Degradation Occurrences (KDO). When a xApp is deployed in the Near-RT-RIC, the MNO is expected to provide the xApp details, including their ICPs ($I_{x_{n}}$) and KPIs ($K_{x_{n}}$). The PGD component of the database is equipped with Algorithm 1, which estimates the $P_{k}^{G}$ for all $k_{j}\in K_{x_{n}}$. The KDO component is responsible for storing and tracking KPI degradation events. If any SLA-sensitive KPI falls below the predefined threshold, it alerts the CDC of a possible conflict, along with the timestamp ($t_{clock}$) indicating when the event occurred.
Figure 4: Average Computation Time for Conflict Detection.
The RCP component tracks any parameter changes, aiding the CDC in identifying the exact parameter responsible for the conflict at $t_{clock}$. The step-by-step conflict detection process is illustrated in Fig. 5, which explains how a rule-based detection mechanism can identify specific types of conflicts. To analyze the performance of the proposed rule-based detection method, we adopted the stochastic xApp model presented in [2, 3] with the same configuration illustrated in Fig. 1. The CDC is triggered only when a KPI violation occurs; otherwise, it is considered a no-conflict state. When the instructing xApp that modifies the conflicting parameter is the same as the xApp associated with the degraded KPI, we classify it as ’no conflict.’ If the instructing xApp and the xApp associated with the degraded KPI are different but share the conflicting parameter, it is detected as a ’direct conflict.’ For an ’indirect conflict,’ we check the parameter group of the degraded KPI; if the conflicting parameter belongs to this group and the case has not been classified as ’no conflict’ or ’direct conflict,’ we consider it an ’indirect conflict.’ Finally, if none of the above three conditions are met, the conflict is classified as ’implicit conflict.’ The rule-based detection method shows an average latency of 0.4 to 0.5 milliseconds for detecting each type of conflict as shown in Fig. 4 with a 100 percent accuracy.
@@
When the conflict is detected, the CDC signals the CMC about the specific types of conflict occurring over the specific conflicting ICP. For the analysis of this paper, we adopt the priority conflict mitigation method proposed by Adamczyk et al. [1] and QACM method proposed in [3] to be deployed in the CMC. The priority method simply priorities one conflicting xApp over the other based on their significance set by the MNO. On the other hand, the QACM considers specific QoS requirements of the conflicting KPIs and suggests an optimal value of the conflicting parameter at which the KPIs are either above their QoS mandates or as closer as possible to it. The following section discusses the simulation setup and results based on these two mitigation methods.
TABLE II: Mobility and Service Type Distribution
Feature Description
Mobility Models Walking (0-1 m/s), Cycling (2-5 m/s), Driving (6-15 m/s)
UE Mobility Distribution (35%) Walking, (30%) Cycling, (35%) Driving
Service Types eMBB (40%), URLLC (30%), mMTC (30%)
Movement Logic Random 2D movement with boundary reflection
V Experiment and Results
For the experimental analysis, we used MATLAB software for simulation, along with its 5G Toolbox and O-RAN 7.2 split [12, 3] and conducted all simulations on a MacBook Pro with an M1 chip and 8GB of RAM. The simulation setup included four Next-generation NodeBs and hundred User Equipments. We focused on a direct conflict between the MRO xApp and ES xApp over Tx Power (TXP) values. The mobility and service type distribution is shown in Table II. The simulation parameters were: a Reference Signal Received Power (RSRP) threshold of -110 dBm for handover, a frequency of 2.4 GHz, and a simulation time of 10 minutes with a time step of 100 ms. Input control parameters included a default TXP value of 30 dBm, Cell Individual Offset (CIO) of 2 dB, Handover Hysteresis (HYS) of 0.5 dB, Time-To-Trigger (TTT) of 0.1 ms, RET of 1.5 degrees, and an adjustment interval of 1000 ms. The UEs moved back and forth between the two gNBs with randomly assigned velocities between 0 and 5 m/s. The overall system goal is to minimize link failures and ping-pong handovers for MRO while maximizing energy efficiency for ES. A direct conflict arose when ES set the TXP to 3 dBm for energy savings, followed by MRO setting it to 50 dBm in the next request. We consider 5 conflict mitigation Methods, including a no-mitigation approach (NC) where the TXP was set to the recently requested value; a set back to default (SBD) method where the TXP was reset to its default value; a priority-based method where either ES (P-ES) or MRO (P-MRO) was prioritized; and finaly the QACM method was used to balance the TXP based on QoS thresholds. We repeated the simulation 500 times and the results were presented using box plots for each KPI.
@@ V Experiment and Results
(c) Number of total handovers.
@@
The Fig. 6(a) compares the energy efficiency across different conflict mitigation methods, including NC (No Conflict Mitigation), SBD (Set Back to Default), P-ES (prioritized ES), P-MRO (prioritized MRO), and QACM. The QACM method shows the highest energy efficiency (in bits/Joule), suggesting it is the most effective approach for maximizing energy savings, followed by P-ES. The NC, SBD and P-MRO methods exhibit lower energy efficiency indicating that prioritizing energy-saving strategies improves this KPI. The QACM optimally sets the transmission power to minimize link failures and unnecessary ping-pong handovers, as reflected in Fig. 6(a). QACM demonstrates 90-100% greater energy efficiency compared to P-ES or P-MRO. This is because none of the prioritized methods effectively balance the transmitted bits-to-energy ratio while achieving their own objectives. SBD and NC exhibit the second-lowest and lowest performance, respectively, as the default values do not optimize any xApp’s objectives. On the other hand, maintaining the value demanded by MRO negatively impacts energy efficiency, since this KPI is not associated with MRO but rather with ES. Therefore, MRO’s instruction doesn’t assure positive impact on energy saving.
@@
The Fig. 6(b) illustrates the impact of different conflict mitigation methods on reducing link failures. The SDB displays higher numbers of link failures followed by the NC and P-ES, with having a similar median value, implying poorer performance in maintaining stable connections. On the other hand, P-MRO and QACM show significantly fewer link failures with QACM being the most effective one. This indicates that prioritizing mobility robustness and a balanced conflict resolution approach reduces the occurrence of link failures. Overall, the QACM reduces approximately 17% of connection losses compared to the NC method.
@@
Finally, the Fig. 6(c) demonstrates the total number of handovers across the various methods. NC, SBD and P-ES have the highest frequency of handovers, which may negatively affect network stability. In contrast, P-MRO and QACM exhibit significantly fewer handovers with QACM providing the best mitigation. This shows that a mobility-optimized and balanced conflict resolution strategy can effectively reduce these undesirable handovers. Overall, the QACM saves approximately 2.2% of unnecessary handovers compared to the NC method.
VI Limitations and Future Works
This study relies on a MATLAB-based simulation with a scenario of one direct conflict between MRO and ES, and a small-scale network of four gNBs and one hundred UEs. This setup may not fully capture real-world complexities such as OTA (Over The Air) interference, multi-cell topologies, or stringent latency demands. In this work, we primarily focused on energy efficiency and did not explore multi-objective or ML-based strategies. This certainly leaves questions about scalability, signaling overhead, and adaptive response of the proposed solutions in rapidly changing environments. Our future work will include deploying on more realistic simulator like ns3-oran [14] developed by NIST, where we have been actively developing and deploying xApps, investigating more diverse conflicts, incorporating multiple objectives, and exploring ML-driven methods for proactive conflict mitigation.
VII Conclusion